# The PLATO field selection process

### v1

@@ 3.1 Main PLATO targets
The histograms of some other astrophysical properties of the P1+P2 and P5 samples are plotted in Fig. 21-22. The spatial distributions of the P1, P2, P4, P5 samples are plotted in the four panels of Fig. 23: P2 and P4 stars are almost perfectly homogeneously distributed, because they both lie relatively close to the Sun, and are purely magnitude-limited samples [^9] Formally, the P2 definition also includes a $\textrm{NSR}\leq 50$ ppm requirement but with the current design it is always satisfied even at its $V=8.5$ faint limit.. P5 is also magnitude-limited, but being on average much fainter is also farther from the Sun, and the consequent line-of-sight effects due to the Galactic disk are revealed by the significant density gradient along Galactic latitude. Finally, P1 is concentrated toward the inner region of LOPS2, observed with 18/24 NCAMs, where the noise requirement is easier to be met at equal magnitude.
Figure 5: Distance of the P1/P2 and P5 samples in LOPS2. Left panel: Distribution of the 8 23582358\,2358 235 P1+P2 stars in LOPS2 in the ($R_{\star}$, distance) plane. The stellar effective temperature from the PIC v2.0.0 is color-coded. Right panel: same, but for the P5 sample (167 149167149167\,149167 149 stars; P1/P2 not included). The distribution moves to larger distances due to the fainter magnitude limit of P5 ($V<13$ vs. $V<11$).
@@ 3.2 Bright stars
Figure 6: Bright stars and stellar crowding in LOPS2. Left panel: orthographic projection in Galactic coordinates centered on LOPS2. The field is color-coded according to the number of NCAMs as in Fig. 1. The brightest stars from the Yale Bright Star Catalog are plotted as circles of different colors: magenta ($V<1$), red ($21<V<$ 2), yellow ($32<V<$ 3), gray ($43<V<$ 4). Right panel: same, but the footprint of LOPS2 is color-coded according to the areal density of $G<13.5$ stars from Gaia DR3. The color scale increases proportionally to density from orange to blue with a transition at $\sim 400$ targets per $\textrm{deg}^{2}$, at $b\simeq-15^{\circ}$.
@@
In any case, knowing the position of such bright sources could be of interest since their bright halos, bleeding columns and smearing trails can potentially contaminate any other fainter target in their neighborhood. A discussion on the technical issues related to CCD saturation, blooming, non-linearity as well as optical ghosting is discussed in Verhoeve et al. (2016); Pertenais et al. (2022); Jannsen et al. (2024) and will be better explored in the forthcoming papers of this series.
@@ 3.3 Nearby stars
Among the 562 known stars within 10 pc from the Sun (Reylé et al. 2021, 2023-02-06 version), 14 are located within LOPS2. Only seven of them are bright enough ($V\lesssim 16$) to be reliably monitored by PLATO [^11] The excluded targets are all at $G\gtrsim 20$, i.e., way too faint for PLATO to deliver reliable photometry.. Six of them are M dwarfs or sub-dwarfs, spanning spectral types from sdM1V to M4.5V; sorted by increasing $G$ magnitude and spectral type: GJ 191, CD-44 3045 B and A, AP Columbae, L 230-188, SCR J0740-4257. The seventh one is LAWD 26 = GJ 293, a bright hydrogen-rich white dwarf (WD; DA; $G\sim 13.7$, $V\sim 14.1$), spectroscopically confirmed by Bell & Rodgers (1964) and recently identified by Sanderson et al. (2022) as one of the most promising white dwarf targets for the discovery of astrometric planets with Gaia.
Other nearby and reasonably bright WDs can be identified from the catalog by Gentile Fusillo et al. (2021): 86 targets with $G<15$ lie within the LOPS2 footprint [^12] We emphasize that WDs will not be included in the main PIC subsamples (P1-P2-P4-P5) due to the spectral class requirement (Section 3.1), unless they are flagged as planetary hosts or candidate hosts (See Section 3.4). Within the techical limits, though, they can be included as “scientific calibration and validation” targets (scvPIC; Zwintz 2024) or as GO targets.. Among these, and besides the already mentioned LAWD 26, four additional stars show a parallax $\pi>50$ mas, i.e., they are closer than 20 pc: two of them were also spectroscopically confirmed by O’Brien et al. (2023), both as DA spectral types.
It is worth mentioning that Kapteyn’s star = GJ 191 ($V\sim 8.8$, $K\sim 5.0$) is an old M1 sub-dwarf at only 3.9 pc from the Sun, and also the nearest known halo star (Kotoneva et al., 2005). It will be observed by PLATO through 12 cameras, and with an average NSR estimated around 18.4 ppm in 1 h according to the current noise budget calculations. Two planetary candidates were claimed to be discovered through RVs around GJ 191 by Anglada-Escude et al. (2014), but were later disproved as due to stellar activity by Robertson et al. (2015) and Bortle et al. (2021), who also noted that the star is photometrically very stable. As will be discussed in Section 3.4, a coordinated RV follow-up of this target could take advantage of the simultaneous PLATO coverage to disentangle the contributions of stellar activity and the Keplerian signal from planets, if any.
@@ 3.4.1 Transiting planets
We retrieved the list of known planets from Exo-MerCat (Alei et al. 2020, Alei et al. 2024 in prep.), a meta-catalog built by comparing and merging with an automated process scanning three input catalogs: Exoplanet Encyclopedia [^13] https://exoplanet.eu/home/, the NASA Exoplanet Archive [^14] https://exoplanetarchive.ipac.caltech.edu/index.html and the Open Exoplanet Catalog [^15] https://www.openexoplanetcatalogue.com/. In this Section we focus on the transiting sub-sample only, which is the most important one given the scientific opportunity they represent for PLATO.
@@
Figure 7: Known planetary systems with at least one transiting planet in LOPS2 (orthographic projection in Galactic coordinates). Left panel: entries from the Exo-MerCat database (Section 3.4): single transiting planets (red circles), systems with multiple transiting planets (orange circles with labels; the number inside the circles is the multiplicity), transiting brown dwarfs (yellow diamonds). Right panel: same projection, but TESS candidates from the TOI database are plotted with yellow circles. Systems with multiple transiting candidates are marked with red squares (doublets) or magenta squares (triplets).
@@
We calculated for each entry of the Exo-MerCat sample the so called ephemeris drift $\sigma(T_{0})$ at epoch 2027.0, i.e., the expected 1-$\sigma$ error on the transit prediction (according to the most accurate ephemeris available) at the approximated epoch when PLATO is supposed to start its scientific operations. The median is $\sigma(T_{0})\sim 7$ min. Only ten planets show a drift larger than two hours; not surprisingly, they are mostly small and/or long-period planets observed by TESS on a limited number of sectors and not accessible by ground-based facilities. At $\sigma(T_{0})\simeq 1.8$ d, an extreme case is TOI-1338b = BEBOP-1b (the first circumbinary planet discovered by TESS, one of the only two known multiple planetary system of this kind; Kostov et al. 2020; Standing et al. 2023), virtually “lost” for the time being. For each of all these loose-ephemeris targets, the first two or three months of PLATO photometry will deliver a new, extremely accurate ephemeris, crucial for the community to setup any further follow-up observations.
@@ 3.4.2 Transiting candidate planets
From the 2024-08-20 release of the TESS Object of Interest database (TOI; Guerrero et al. 2021) we identified 824 entries in LOPS2 and matched them with the TESS Follow-up program (TFOP) data base, which includes a disposition flag (“TFOPWG”) reflecting the current status of each target according to the follow-up results. After selecting only successfully vetted planetary candidates (disposition keyword: PC, APC or CP, meaning “planetary candidate”, “ambiguous planetary candidate” and “confirmed planet”, respectively) not already published we are left with 544 candidates hosted by 513 stars [^18] The full list will be made available online on Zenodo. (Fig. 7, right panel). We emphasize that these numbers include candidates with a very wide range of FPR, and that a quick inspection at the “comment” column of the TOI database reveals that a considerable fraction of APC entries could likely turn out as false positives. Nevertheless, an individual vetting of each TOI is beyond the scope of this paper. Notably, 27 stars among the LOPS2 TOI host multiple candidates, a configuration which is known to significantly decrease the a-priori FPR. Three systems have multiplicity $N=3$: TOI-699, TOI-790, and TOI-2392 currently under active follow-up by the TFOP working group.
It is worth mentioning that 27 among our list of TESS candidates are listed in the TOI table as having orbital periods longer than 100 d (they are visible at the right end of Fig. 8, right panel) or with no period at all. A closer inspection at the comment field reveals that almost all of them are so called “monotransits”, i.e., candidates detected through a single event whose orbital period is unconstrained. In fact, an even larger number of TOI candidates, especially in the long-period end, are flagged as having an ambiguous period due to an insufficient orbital phase coverage by TESS. The almost uninterrupted 2-year coverage of LOPS2 will easily break the ambiguity and detect the correct orbital period (Magliano et al., 2024).
3.4.3 Non transiting planets
@@
  • Radial velocity: 64 planets grouped in 41 planetary systems. Among the 15 multiple systems, HD 40307 (Mayor et al., 2009) stands out as that with the highest multiplicity: at least four planets confirmed by independent analyses, plus two more controversial candidates [^19] As a side note, the inclusion of HD 40307 and other bright RV systems in LOPS2 brings up the opportunity of planning a $\geq 2$ yr observing campaign where ultra-high-precision RV measurements can be combined with a simultaneous and almost uninterrupted space-based light curve to disentangle the planetary signal from stellar activity effects, following the approach developed by Lanza et al. (2011); Aigrain et al. (2012); Haywood et al. (2014), among others. (Tuomi et al., 2013; Díaz et al., 2016). A few RV systems are not counted here because they also host confirmed transiting planets, so they were included in the sample described in Section  3.4.1: they are TOI-500 and TOI-431 (both hosting a transiting super-Earth), L98-59 (with three transiting planets, Kostov et al. 2019; Demangeon et al. 2021), and TOI-1338A = BEBOP-1 (Standing et al., 2023);
  • Direct imaging: 11 planets hosted by ten systems, among which $\beta$ Pic b (Lagrange et al., 2010, 2019), AB Pic (Chauvin et al., 2005), HR 2562 (Konopacky et al., 2016), and WD 0806-661 b (Luhman et al., 2011) stand out as the best characterized. All of them are very massive planets, close to or across the planet-BD transition;
@@
Figure 9: Known non-transiting planetary systems in LOPS2 (orthographic projection in Galactic coordinates). Left panel: entries from the Exo-MerCat database (Section 3.4): planets discovered through RVs (red diamonds), direct imaging (yellow circles), astrometry (gray circle), and ETV (green circle). The location of the candidate TTV system WASP-126 is also marked with a magenta point. Right panel: Candidate astrometric planets published by the Gaia collaboration (purple circles); each data point is labelled with its ASOI ID number.
@@ 3.5 Star clusters and associations
Figure 10: Open clusters and associations in LOPS2. Left panel: sky map of the 367 clusters identified by Hunt & Reffert (2023) at high confidence and with at least one member in LOPS2, plotted as red circles with radius r50 (radius containing 50% of the members within the tidal radius). All the 10 682 members at $V<15$ are plotted as black points. Right panel: The eight clusters in our sample having the largest number of $V<15$ targets of spectral type F5V. Each target is color-coded with the same scheme as in Fig. 11.
Figure 11: Open clusters and associations in LOPS2. Left panel: the 367 clusters identified by Hunt & Reffert (2023) at high confidence and with at least one member in LOPS2, plotted as a function of their distance and areal density (as defined in the text). The logarithmic age of each cluster is color-coded, while the point area is proportional to the number of $V<15$ stars. The eight clusters having the largest number of $V<15$ targets of spectral type F5V and later are labelled. Right panel: absolute and de-reddened color-magnitude diagram (CMD) for all the 10 6821068210\,68210 682 $V<15$ stars in LOPS2 belonging to clusters (gray points). Members of the eight labeled clusters are plotted with larger circles and color-coded as in the legend.
@@ 3.6 Variable and binary stars
Thanks to its very large area (2132 $\textrm{deg}^{2}$, i.e., 5.2% of the whole sky) and to the wide range in Galactic latitude covered, LOPS2 will include a large number and variety of variable stars. A cross-match with the latest versions of the VSX (Variable Star indeX; Watson et al. 2006) and Gavras et al. (2023) catalogs yields 47 3564735647\,35647 356 and 282 366282366282\,366282 366 entries, respectively (with notable over-densities in the LMC and close to the Galactic plane). The numbers decrease to 31 2113121131\,21131 211 and 15 9321593215\,93215 932, respectively, when we limit our sample at $V<15$. Among these we mention:
@@
  • 185 chemically-peculiar stars;
  • 47 cataclismic variables (CV);
@@
  • 29 slowly-pulsating B stars (SPB);
@@
It should be emphasized that some of these classes are interesting not just for stellar science, but also to plan more focused planet-hunting surveys. For instance, circumbinary planets represent a rare laboratory to challenge our theories on planetary formation and migration through a disk, and can be discovered both by detecting their transits (Kostov et al., 2020) [^20] The first (and so far only) circumbinary planet detected by TESS, TOI-1338 (Kostov et al., 2020), is by chance in LOPS2., or by modeling their eclipse timing variations (ETV; Goldberg et al. 2023; Brown-Sevilla et al. 2021). Pulsators with intrinsically stable modes, such as $\delta$ Scuti and $\gamma$ Dor variables, may enable the detection of non-transiting planets on wide orbits through the pulsation timing technique, thanks to the unique combination of timing accuracy and temporal baseline of PLATO (Vaulato et al., 2022). Early-type stars with very specific chemical signatures such as $\lambda$ Bootis stars have been predicted since long to host a much higher fraction of giant planets with respect to ordinary field stars, although this has been recently disputed (Saffe et al., 2021).
@@
  • R Doradus, an asymptotic giant branch (AGB) star with a 300 $R_{\odot}$ radius, probably the star with the largest apparent diameter as seen from the Sun.
Aside from these few very bright variables and the cross-match with the catalog by Gavras et al. (2023), we have also searched for confirmed non-radial pulsators in LOPS2. This was done from a combined Gaia–TESS approach. Hey & Aerts (2024) distilled about 60 0006000060\,00060 000 variables using light curves from the first year of the TESS mission, starting from the original Gaia DR3 catalog of candidate main-sequence pulsators by Gaia Collaboration et al. (2023b) relying on their stellar properties derived by Aerts et al. (2023). Hey & Aerts (2024) reclassified all the stars whose dominant frequency in the totally independent sparsely sampled Gaia DR3 light curves and the 30-min sampled TESS light curves are equal. All these stars have a dominant amplitude above 4 mmag in the Gaia $G$ band. Among the 6430 stars from Hey & Aerts (2024) occurring in LOPS2, about 5000 are now classified as confirmed multi-periodic pulsators, while the others are rotational variables or eclipsing binaries. These non-radial pulsators in LOPS2 are split up into 1455 gravity-mode pulsators (of $\gamma\,$Dor or SPB type), 2079 $\delta\,$Sct pulsators, and 1449 hybrid pressure- and gravity mode pulsators. Mombarg et al. (2024) provide estimates of the (convective core) masses, radii, and relative ages for more than 10 000 of the $\gamma\,$Dor, SPB, and hybrid pulsators classified by Hey & Aerts (2024). They did so by subjecting the pulsators’ Gaia data to grid-based modelling based on rotating stellar models. These models were constructed such as to be compliant with the asteroseismology of rotating Kepler pulsators in the mass range from 1.3 $M_{\odot}$ and 9 $M_{\odot}$ from Li et al. (2020); Pedersen et al. (2021). Some 3000 of the stars treated by Mombarg et al. (2024) occur in LOPS2. For about 300 among them Aerts et al. (2024) distilled the near-core rotation frequency from their identified dominant dipole prograde gravito-inertial mode as found consistently in both the Gaia DR3 and TESS light curves. Ongoing work on the full 5-year TESS light curves will increase the number of new pulsators with a measurement of the internal rotation frequency and this estimate will become more precise. Moreover, simulations with PLATOSIM (Jannsen et al., 2024) reveal that detections of a dense spectrum of oscillation modes suitable for asteroseismology are expected for all these $\gamma\,$Dor and hybrid pulsators brighter than $V<14$ (Jannsen et al., in prep.), making them ideal calibration stars for the stellar science program of the mission.

### v3

@@ 3.1 Main PLATO targets
The histograms of some other astrophysical properties of the P1+P2 and P5 samples are plotted in Fig. 21-22. The spatial distributions of the P1, P2, P4, P5 samples are plotted in the four panels of Fig. 23: P2 and P4 stars are almost perfectly homogeneously distributed, because they both lie relatively close to the Sun, and are purely magnitude-limited samples. [^9] Formally, the P2 definition also includes a $\textrm{NSR}\leq 50$ ppm requirement but with the current design it is always satisfied even at its $V=8.5$ faint limit. P5 is also magnitude-limited, but being on average much fainter is also farther from the Sun, and the consequent line-of-sight effects due to the galactic disk are revealed by the significant density gradient along galactic latitude. Finally, P1 is concentrated toward the inner region of LOPS2, observed with 18/24 NCAMs, where the noise requirement is easier to be met at equal magnitude.
Figure 5: Distance of the P1/P2 and P5 samples in LOPS2. Left panel: Distribution of the 8 23582358\,2358 235 P1+P2 stars in LOPS2 in the ($R_{\star}$, distance) plane. The stellar effective temperature from the PIC v2.0.0 is color coded. Right panel: Same but for the P5 sample (167 149167149167\,149167 149 stars; P1/P2 not included). The distribution moves to larger distances due to the fainter magnitude limit of P5 ($V<13$ vs. $V<11$).
@@ 3.2 Bright stars
Figure 6: Bright stars and stellar crowding in LOPS2. Left panel: Orthographic projection in galactic coordinates centered on LOPS2. The field is color coded according to the number of NCAMs as in Fig. 1. The brightest stars from the Yale Bright Star Catalog are plotted as circles of different colors: magenta ($V<1$), red ($21<V<$ 2), yellow ($32<V<$ 3), gray ($43<V<$ 4). Right panel: Same but the footprint of LOPS2 is color coded according to the areal density of $G<13.5$ stars from Gaia DR3. The color scale increases proportionally to density from orange to blue with a transition at $\sim 400$ targets per $\textrm{deg}^{2}$, at $b\simeq-15^{\circ}$.
@@
In any case, knowing the position of such bright sources could be of interest since their bright halos, bleeding columns and smearing trails can potentially contaminate any other fainter target in their neighborhood. A discussion on the technical issues related to CCD saturation, blooming, non-linearity as well as optical ghosting is discussed in Verhoeve et al. (2016); Pertenais et al. (2022); Jannsen et al. (2024a) and will be better explored in the forthcoming papers of this series.
@@ 3.3 Nearby stars
Among the 562 known stars within 10 pc from the Sun (Reylé et al. 2021, 2023-02-06 version), 14 are located within LOPS2. Only seven of them are bright enough ($V\lesssim 16$) to be reliably monitored by PLATO. [^11] The excluded targets are all at $G\gtrsim 20$, that is, way too faint for PLATO to deliver reliable photometry. Six of them are M dwarfs or sub-dwarfs, spanning spectral types from sdM1V to M4.5V; sorted by increasing $G$ magnitude and spectral type: GJ 191, CD-44 3045 B and A, AP Columbae, L 230-188, SCR J0740-4257. The seventh one is LAWD 26 = GJ 293, a bright hydrogen-rich white dwarf (WD; DA; $G\sim 13.7$, $V\sim 14.1$), spectroscopically confirmed by Bell & Rodgers (1964) and recently identified by Sanderson et al. (2022) as one of the most promising white dwarf targets for the discovery of astrometric planets with Gaia.
Other nearby and reasonably bright WDs can be identified from the catalog by Gentile Fusillo et al. (2021): 86 targets with $G<15$ lie within the LOPS2 footprint. [^12] We emphasize that WDs will not be included in the main PIC subsamples (P1-P2-P4-P5) due to the spectral class requirement (Section 3.1), unless they are flagged as planetary hosts or candidate hosts (See Section 3.4). Within the technical limits, though, they can be included as “scientific calibration and validation” targets (scvPIC; Zwintz 2024) or as GO targets. Among these, and besides the already mentioned LAWD 26, four additional stars show a parallax $\pi>50$ mas, that is, they are closer than 20 pc: two of them were also spectroscopically confirmed by O’Brien et al. (2023), both as DA spectral types.
It is worth mentioning that Kapteyn’s star = GJ 191 ($V\sim 8.8$, $K\sim 5.0$) is an old M1 sub-dwarf at only 3.9 pc from the Sun, and also the nearest known halo star (Kotoneva et al., 2005). It will be observed by PLATO through 12 cameras, and with an average NSR estimated around 18.4 ppm in 1 h according to the current noise budget calculations. Two planetary candidates were claimed to be discovered through RVs around GJ 191 by Anglada-Escude et al. (2014), but were later disproved as due to stellar activity by Robertson et al. (2015) and Bortle et al. (2021), who also noted that the star is photometrically very stable. As is discussed in Section 3.4, a coordinated RV follow-up of this target could take advantage of the simultaneous PLATO coverage to disentangle the contributions of stellar activity and the Keplerian signal from planets, if any.
@@ 3.4.1 Transiting planets
We retrieved the list of known planets from Exo-MerCat (Alei et al. 2020, Alei et al. in prep.), a meta-catalog built by comparing and merging with an automated process scanning three input catalogs: Exoplanet Encyclopedia, [^13] https://exoplanet.eu/home/ the NASA Exoplanet Archive [^14] https://exoplanetarchive.ipac.caltech.edu/index.html and the Open Exoplanet Catalog. [^15] https://www.openexoplanetcatalogue.com/ In this Section we focus on the transiting sub-sample only, which is the most important one given the scientific opportunity they represent for PLATO.
@@
Figure 7: Known planetary systems with at least one transiting planet in LOPS2 (orthographic projection in galactic coordinates). Left panel: Entries from the Exo-MerCat database (Section 3.4): single transiting planets (red circles), systems with multiple transiting planets (orange circles with labels; the number inside the circles is the multiplicity), transiting brown dwarfs (yellow diamonds). Right panel: Same projection but TESS candidates from the TOI database are plotted with yellow circles. Systems with multiple transiting candidates are marked with red squares (doublets) or magenta squares (triplets).
@@
We calculated for each entry of the Exo-MerCat sample the so called ephemeris drift $\sigma(T_{0})$ at epoch 2027.0, that is, the expected 1-$\sigma$ error on the transit prediction (according to the most accurate ephemeris available) at the approximated epoch when PLATO is supposed to start its scientific operations. The median is $\sigma(T_{0})\sim 7$ min. Only ten planets show a drift larger than two hours; not surprisingly, they are mostly small and/or long-period planets observed by TESS on a limited number of sectors and not accessible by ground-based facilities. At $\sigma(T_{0})\simeq 1.8$ d, an extreme case is TOI-1338b = BEBOP-1b (the first circumbinary planet discovered by TESS, one of the only two known multiple planetary system of this kind; Kostov et al. 2020; Standing et al. 2023), virtually “lost” for the time being. For each of all these loose-ephemeris targets, the first two or three months of PLATO photometry will deliver a new, extremely accurate ephemeris, crucial for the community to setup any further follow-up observations.
@@ 3.4.2 Transiting candidate planets
From the 2024-08-20 release of the TESS Object of Interest database (TOI; Guerrero et al. 2021) we identified 824 entries in LOPS2 and matched them with the TESS Follow-up program (TFOP) data base, which includes a disposition flag (“TFOPWG”) reflecting the current status of each target according to the follow-up results. After selecting only successfully vetted planetary candidates (disposition keyword: PC, APC or CP, meaning “planetary candidate”, “ambiguous planetary candidate” and “confirmed planet”, respectively) not already published we are left with 544 candidates hosted by 513 stars [^18] The full list is available on zenodo14720127. (Fig. 7, right panel). We emphasize that these numbers include candidates with a very wide range of FPR, and that a quick inspection at the “comment” column of the TOI database reveals that a considerable fraction of APC entries could likely turn out as false positives. Nevertheless, an individual vetting of each TOI is beyond the scope of this paper. Notably, 27 stars among the LOPS2 TOI host multiple candidates, a configuration which is known to significantly decrease the a-priori FPR. Three systems have multiplicity $N=3$: TOI-699, TOI-790, and TOI-2392 currently under active follow-up by the TFOP working group.
It is worth mentioning that 27 among our list of TESS candidates are listed in the TOI table as having orbital periods longer than 100 d (they are visible at the right end of Fig. 8, right panel) or with no period at all. A closer inspection at the comment field reveals that almost all of them are so called “monotransits”, that is, candidates detected through a single event whose orbital period is unconstrained. In fact, an even larger number of TOI candidates, especially in the long-period end, are flagged as having an ambiguous period due to an insufficient orbital phase coverage by TESS. The almost uninterrupted 2-year coverage of LOPS2 will easily break the ambiguity and detect the correct orbital period (Magliano et al., 2024).
3.4.3 Non-transiting planets
@@
  • Radial velocity: sixty-four planets grouped in 41 planetary systems. Among the 15 multiple systems, HD 40307 (Mayor et al., 2009) stands out as that with the highest multiplicity: at least four planets confirmed by independent analyses, plus two more controversial candidates [^19] As a side note, the inclusion of HD 40307 and other bright RV systems in LOPS2 brings up the opportunity of planning a $\geq 2$ yr observing campaign where ultra-high-precision RV measurements can be combined with a simultaneous and almost uninterrupted space-based light curve to disentangle the planetary signal from stellar activity effects, following the approach developed by Lanza et al. (2011); Aigrain et al. (2012); Haywood et al. (2014), among others. (Tuomi et al., 2013; Díaz et al., 2016). A few RV systems are not counted here because they also host confirmed transiting planets, so they were included in the sample described in Section  3.4.1: they are TOI-500 and TOI-431 (both hosting a transiting super-Earth), L98-59 (with three transiting planets, Kostov et al. 2019; Demangeon et al. 2021), and TOI-1338A = BEBOP-1 (Standing et al., 2023);
  • Direct imaging: eleven planets hosted by ten systems, among which $\beta$ Pic b (Lagrange et al., 2010, 2019), AB Pic (Chauvin et al., 2005), HR 2562 (Konopacky et al., 2016), and WD 0806-661 b (Luhman et al., 2011) stand out as the best characterized. All of them are very massive planets, close to or across the planet-BD transition;
@@
Figure 9: Known non-transiting planetary systems in LOPS2 (orthographic projection in galactic coordinates). Left panel: Entries from the Exo-MerCat database (Section 3.4): planets discovered through RVs (red diamonds), direct imaging (yellow circles), astrometry (gray circle), and ETV (green circle). The location of the candidate TTV system WASP-126 is also marked with a magenta point. Right panel: Candidate astrometric planets published by the Gaia collaboration (purple circles). Each data point is labelled with its ASOI ID number.
@@ 3.5 Star clusters and associations
Figure 10: Open clusters and associations in LOPS2. Left panel: Sky map of the 367 clusters identified by Hunt & Reffert (2023) at high confidence and with at least one member in LOPS2, plotted as red circles with radius r50 (radius containing 50% of the members within the tidal radius). All the 10 682 members at $V<15$ are plotted as black points. Right panel: Eight clusters in our sample having the largest number of $V<15$ targets of spectral type F5V. Each target is color coded with the same scheme as in Fig. 11.
Figure 11: Open clusters and associations in LOPS2. Left panel: the 367 clusters identified by Hunt & Reffert (2023) at high confidence and with at least one member in LOPS2, plotted as a function of their distance and areal density (as defined in the text). The logarithmic age of each cluster is color coded, while the point area is proportional to the number of $V<15$ stars. The eight clusters having the largest number of $V<15$ targets of spectral type F5V and later are labelled. Right panel: Absolute and de-reddened color-magnitude diagram (CMD) for all the 10 6821068210\,68210 682 $V<15$ stars in LOPS2 belonging to clusters (gray points). Members of the eight labeled clusters are plotted with larger circles and color coded as in the legend.
@@ 3.6 Variable and binary stars
Thanks to its very large area (2149 $\textrm{deg}^{2}$, that is, 5.2% of the whole sky) and to the wide range in galactic latitude covered, LOPS2 will include a large number and variety of variable stars. A cross-match with the latest versions of the VSX (Variable Star indeX; Watson et al. 2006) and Gavras et al. (2023) catalogs yields 47 3564735647\,35647 356 and 282 366282366282\,366282 366 entries, respectively (with notable over-densities in the LMC and close to the galactic plane). The numbers decrease to 31 2113121131\,21131 211 and 15 9321593215\,93215 932, respectively, when we limit our sample at $V<15$. Among these we mention:
@@
  • 185 chemically peculiar stars;
  • 47 cataclysmic variables (CV);
@@
  • 29 slowly pulsating B stars (SPB);
@@
It should be emphasized that some of these classes are interesting not just for stellar science, but also to plan more focused planet-hunting surveys. For instance, circumbinary planets represent a rare laboratory to challenge our theories on planetary formation and migration through a disk, and can be discovered both by detecting their transits (Kostov et al., 2020), [^20] The first (and so far only) circumbinary planet detected by TESS, TOI-1338 (Kostov et al., 2020), is by chance in LOPS2. or by modeling their eclipse timing variations (ETV; Goldberg et al. 2023; Brown-Sevilla et al. 2021). Pulsators with intrinsically stable modes, such as $\delta$ Scuti and $\gamma$ Dor variables, may enable the detection of non-transiting planets on wide orbits through the pulsation timing technique, thanks to the unique combination of timing accuracy and temporal baseline of PLATO (Vaulato et al., 2022). Early-type stars with very specific chemical signatures such as $\lambda$ Bootis stars have been predicted since long to host a much higher fraction of giant planets with respect to ordinary field stars, although this has been recently disputed (Saffe et al., 2021).
@@
  • R Doradus, an asymptotic giant branch (AGB) star with a 300 $R_{\odot}$ radius, probably the star with the largest apparent diameter as seen from the Sun (Bedding et al., 1997).
Aside from these few very bright variables and the cross-match with the catalog by Gavras et al. (2023), we have also searched for confirmed non-radial pulsators in LOPS2. This was done from a combined Gaia–TESS approach. Hey & Aerts (2024) distilled about 60 0006000060\,00060 000 variables using light curves from the first year of the TESS mission, starting from the original Gaia DR3 catalog of candidate main-sequence pulsators by Gaia Collaboration et al. (2023b) relying on their stellar properties derived by Aerts et al. (2023). Hey & Aerts (2024) reclassified all the stars whose dominant frequency in the totally independent sparsely sampled Gaia DR3 light curves and the 30-min sampled TESS light curves are equal. All these stars have a dominant amplitude above 4 mmag in the Gaia $G$ band. Among the 6430 stars from Hey & Aerts (2024) occurring in LOPS2, about 5000 are now classified as confirmed multi-periodic pulsators, while the others are rotational variables or eclipsing binaries. These non-radial pulsators in LOPS2 are split up into 1455 gravity-mode pulsators (of $\gamma\,$Dor or SPB type), 2079 $\delta\,$Sct pulsators, and 1449 hybrid pressure- and gravity mode pulsators. Mombarg et al. (2024) provide estimates of the (convective core) masses, radii, and relative ages for more than 10 000 of the $\gamma\,$Dor, SPB, and hybrid pulsators classified by Hey & Aerts (2024). They did so by subjecting the pulsators’ Gaia data to grid-based modelling based on rotating stellar models. These models were constructed such as to be compliant with the asteroseismology of rotating Kepler pulsators in the mass range from 1.3 $M_{\odot}$ and 9 $M_{\odot}$ from Li et al. (2020); Pedersen et al. (2021). Some 3000 of the stars treated by Mombarg et al. (2024) occur in LOPS2. For about 300 among them Aerts et al. (A&A subm.) distilled the near-core rotation frequency from their identified dominant dipole prograde gravito-inertial mode as found consistently in both the Gaia DR3 and TESS light curves. Ongoing work on the full 5-year TESS light curves will increase the number of new pulsators with a measurement of the internal rotation frequency and this estimate will become more precise. Moreover, simulations with PLATOSIM (Jannsen et al., 2024a) reveal that detections of a dense spectrum of oscillation modes suitable for asteroseismology are expected for all these $\gamma\,$Dor and hybrid pulsators brighter than $V<14$ (Jannsen et al., 2024b), making them ideal calibration stars for the stellar science program of the mission.